OpenStudy (anonymous):

at 7:30 how many degrees are included between the hand of a clock

OpenStudy (anonymous):

7 is at7/12 of the whole circle = (7/12)*360 = 210 and 30 is 6/12 = 1/2 of circle which is 36-x 1/2 = 180 so difference 210-180=30